Hallux abducto valgus (call it HAV) is a deformation of the big toe that often requires surgery. Doctors used X-rays to measure the angle (in degrees) of deformity in 38 consecutive patients under the age of 21 who came to a a medical center for surgery to correct HAV, The angle is a measure of the seriousness of the deformity. Here are the data. 29 33 25 34 39 24 25 16 28 28 26 11 18 19 23 21 13 34 24 21 21 22 18 26 16 31 32 20 52 24 27 29 32 40 34 21 It is reasonable to regard these patients as a random sample of young patients who require HAV surgery. Carry out the Solve and Conclude steps of a 95% confidence interval for the mean HAV angle in the population of all such patients. (Round your answer for x to two decimal places and your answer for s to three decimal places. Round your confidence interval bounds to two decimal places.) confidence interval bounds
Answer #1

From the given data, the following statistics are calculated:

n = Sample size = 38

\bar{x} = Sample mean = 976/38 = 25.6842

s = Sample SD = 8.1177

SE = s/\sqrt{n} = 8.1177/V38 = 1.3169

\alpha = 0.05

From Table, Critical value of Z = 1.96.

95% Confidence Interval:

\bar{x}\pm Z SE

= 25.6842 \pm 1.96 X 1.3169

= 25.6842 \pm 2.5810

=( 23.10, 28.27)

So, the required answers are:

\bar{x} = 25.68

s = 8.118

confidence interval bounds:

23.10 to 28.27